MIDC News
Year of 2018
October 29, 2018
- 09:17 MST: New MIDC Feature: The x and y axis of site comparison plots can now be rescaled. -Afshin Andreas
- 14:08 MST: August through September 2018 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 07:23 MST: The NWTC M2 tower was dropped on Oct 12 from 08:07 to 11:51 MST, so that the temperature probes, wind speed and direction sensors could replaced. Calibration constants for new temperature and wind speed sensors were updated in logger program at 09:52 MST on Oct 12. -Afshin Andreas
- 14:38 MST: A new measurement station at the University of Louisiana at Lafayette has been added to MIDC. Data is available since August 28, 2018. -Afshin Andreas
- 15:39 MST: The University of Arizona (OASIS) has been down since September 2 due to a problem with the data logger's communication interface. -Afshin Andreas
- 13:20 MST: The SRRL BMS Upwelling Quantum LI-190 sensor was calibrated today. New factor was entered in logger at 13:16 MST today. -Afshin Andreas
- 19:18 MST: April through July 2018 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 11:45 MST: A new 1-Axis MS-711 Spectroradiometer was installed at the University of Oregon. -Afshin Andreas
- 11:26 MST: A new Dew Point transducer was installed at NWTC M2 at 10:50 MST today. -Afshin Andreas
- 17:23 MST: NWTC M2 collection was restarted at approx. 13:00 MST. A new dewpoint transducer is yet to be installed, therefore the dewpoint and relative humidity readings are bogus. In addition, the 80m temperature is not operational at this time. -Afshin Andreas
- 09:45 MST: Data collection for NWTC M2 was turned off at 09:23 MST today, so that the data logger can be upgraded from a CR23X to a CR3000. -Afshin Andreas
- 07:39 MST: The SRRL BMS UV-S-A-T power/signal cable was replaced on 8/17/2018 from 11:15 to 11:22 MST. There is no data from the UV-S-A-T & UV-S-B-T during that time. The cable had suspected internal damage from the hail storm in May 2017. -Afshin Andreas
- 14:25 MST: The SRRL BMS Deck and Radiometer Tower CR3000 data loggers were calibrated today. No measurements were affected. -Afshin Andreas
- 08:51 MST: SRRL BMS Global 40-South LI-200 sensor was temporarily swapped with a calibrated spare until the original sensor can be repaired. -Afshin Andreas
- 07:59 MST: The was a large rain/hail storm at SRRL yesterday @ 16:16 MST. The following problems have occurred as a result: Global 40-South LI-200 is reading zero. Global 90-North LI-200 and Global UV-S-A-T have developed a positive offset. -Afshin Andreas
- 10:44 MST: At 10:43 MST today, a Global 40-South Total-UV CUV4 was added to SRRL BMS. Instrument is mounted next to Global 40-South CMP11 on radiometer tower. -Afshin Andreas
- 15:03 MST: The UCAR web site is operational again. PWV data for SRRL, University of Oregon, and SolarTAC are now up to date on MIDC. -Afshin Andreas
- 12:17 MST: The SRRL BMS Global CUV4 was replaced with another calibrated CUV4. The original CUV4 will become the future SRRL BMS Global 40-South CUV4. Logger was updated with new factor @ 12:10 MST today. -Afshin Andreas
- 19:42 MST: As of 08:40 MST on 8/2/2018, the SRRL BMS UV-S-A-T & UV S-B-T include a voltage offset to remove the offset created by the control electronics. -Afshin Andreas
- 19:39 MST: All SRRL BMS UV Radiometers (global and direct) were calibrated on 7/31/2018. New factors were entered into logger on 8/2/2018 @ 07:36 MST. -Afshin Andreas
- 13:50 MST: The SRRL BMS UVB-1 (which was a spare) was swapped back to the primary unit today @ 12:15 MST. -Afshin Andreas
- 09:08 MST: The SRRL BMS Global Upwelling CMP11 and Diffuse Shadow Band 8-48 were swapped out with instruments calibrated in BORCAL SW 2018-04 and 2018-05 today. Logger was updated with new factors @ 08:53 MST today. -Afshin Andreas
- 09:06 MST: The SRRL BMS Global Upwelling CM3 and LI200 were swapped out with instruments calibrated in BORCAL SW 2018-04 today. Logger was updated with new factors @ 08:08 MST today. -Afshin Andreas
- 09:06 MST: The SRRL BMS Global 40-South CMP11 and LI200 were swapped out with instruments calibrated in BORCAL SW 2018-04 today. Logger was updated with new factors @ 07:43 MST today. -Afshin Andreas
- 07:25 MST: The SRRL BMS Global RG780 PSP and Photometric LI-210 sensors were calibrated on 05/16/2018. New factor was entered in the logger @ 07:18 MST today. -Afshin Andreas
- 14:41 MST: The SRRL BMS Diffuse CM22 #2 was replaced with an instrument calibrated in BORCAL SW 2018-04 at 14:37 MST today. -Afshin Andreas
- 12:33 MST: The SRRL BMS Global PQS1 PAR sensor was calibrated on 5/16/2018. New factor was entered in logger @ 11:31 MST today. Also at that time the LI-190 was replaced with another sensor (the old sensor would get water contaminated by rain then dry out over time). -Afshin Andreas
- 12:22 MST: SRRL BMS Pressure sensor was swapped out with a calibration sensor at 12:15 MST today. -Afshin Andreas
- 12:53 MST: Per DOE Requirements http has been disabled on the public MIDC web server. Web browsers should automatically redirect to https. Users that grab data programmatically may have to update their software to use port 443 instead of port 80. -Afshin Andreas
- 14:36 MST: The UCAR web site has been down since June 12. Since they process PWV data for NREL there is no updated PWV data for SRRL, University of Oregon, and SolarTAC since then. -Afshin Andreas
- 15:55 MST: All SRRL BMS Direct Normal and Global Horizontal Shortwave Radiometers were calibrated in BORCAL SW 2018-03 (via signal patching), this includes: CMP22 (vent), CMP22 #1, CMP22 #2, CMP11, CM6b, CM3, SR25, MS-80, MS-802, MS-410, MS-602, SPP, PSP, GPP, SP-510, TSP-700, SP-Lite2, ML-01, LI-200R, LI-200, SP-110; CHP1 #1, CHP1 #2, sNIP, NIP, DR02, MS-56, LI-201. Logger was updated with new factors at 15:23 MST today (15:25 MST for PSP RSvZ). -Afshin Andreas
- 15:49 MST: The SRRL BMS Global Normal CMP22 and LI-200R were re-installed at 10:54 MST on May 24 using sensors that were calibrated in BORCAL SW 2018-03. -Afshin Andreas
- 16:30 MST: The SRRL BMS Direct MS-56 has a large positive offset that started on May 23 @ 8:00 MST. Problem will be investigated tomorrow. -Afshin Andreas
- 10:53 MST: The SRRL BMS Global Normal CMP22 and LI-200R were removed by mistake at 10:11 MST today. They will been re-installed soon. -Afshin Andreas
- 16:09 MST: The SRRL BMS Diffuse SR25 and 8-48 were replaced with an instrument calibrated in BORCAL SW 2018-03 at 15:56 MST today. -Afshin Andreas
- 15:15 MST: The SRRL BMS Diffuse CM22 #1 was replaced with an instrument calibrated in BORCAL SW 2018-03 at 14:28 MST today. -Afshin Andreas
- 11:19 MST: The following SRRL BMS Direct Normal instruments were calibrated on 5/8/2018: PGS-100, LI-190, and NIP RG780. Factors for LI-190 and NIP RG780 were updated in logger at 11:12 MST. PGS-100 was well within tolerance and no adjustment was made to calibration file. -Afshin Andreas
- 10:56 MST: The SRRL BMS Global Quantum LI-190 has been reading low since May 14 due to water contamination. The PQS-1 should be used until sensor dries (or is replaced). -Afshin Andreas
- 06:25 MST: A GFCI tripped or failed, on the SRRL instrument deck near midnight. The following SRRL BMS instruments have no power and will read missing or bogus: Direct UVA, Direct UVB, PGS-100, POM-01, TWC. -Afshin Andreas
- 12:14 MST: The SRRL BMS 90-North PSP and LI-200 were swapped out with sensors calibrated in BORCAL 2017-02 between 11:25 and 11:40 MST today. Logger was updated with new calibration factors at 11:25 MST. -Afshin Andreas
- 12:12 MST: The SRRL BMS 90-South PSP and LI-200 were swapped out with sensors calibrated in BORCAL 2017-02 between 10:48 and 11:04 MST today. Logger was updated with new calibration factors at 10:49 MST. -Afshin Andreas
- 15:15 MST: The SRRL EKO ASI-16 Sky Camera is currently down, there will be no updated images or data until the firmware can be updated/repaired. -Afshin Andreas
- 15:28 MST: January through March 2018 Monthly Data Sets & Reports/Plots for SRRL BMS were generated yesterday. -Afshin Andreas
- 09:04 MST: Due to upcoming BORCAL logistics, the BMS Diffuse SR25 had to be swapped with the spare SR25 (which was calibrated in BORCAL 2017-01) at 08:08 MST today. -Afshin Andreas
- 09:45 MST: The SRRL BMS TSP-700 has been reading low since April 1 (up to about 100 W/m^2) at noon, problem unknown at this time. -Afshin Andreas
- 14:43 MST: The Sun Edison solar trackers on the mesa top failed on March 17 at 15:00 MST. The BMS SE-POA LI-200 data should not be used until the solar trackers are repaired. -Afshin Andreas
- 11:54 MST: The OTF MS-700 Spectroradiometer was removed from service at 11:24 MST on Feb 15 for calibration. However due to the large loss in response it was sent to the factory for possible repair. -Afshin Andreas
- 10:11 MST: The SRRL BMS 90-West PSP and LI-200 were swapped out with sensors calibrated in BORCAL 2017-02 between 9:20 and 9:40 MST today. Logger was updated with new calibration factors at 9:27 MST. -Afshin Andreas
- 10:10 MST: The SRRL BMS 90-East PSP and LI-200 were swapped out with sensors calibrated in BORCAL 2017-02 between 8:50 and 9:10 MST today. Logger was updated with new calibration factors at 9:03 MST. -Afshin Andreas
- 15:37 MST: An EKO All Sky Imager Model ASI-16 was installed at SRRL on 9/26/2017 to replace the NREL-built SkyCam that was damaged by the May 2017 hail storm. Images from the camera are now available on MIDC. The post processing software uses two algorithms to produce cloud cover percentages and post processed images. -Afshin Andreas
- 11:38 MST: December 2017 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 15:54 MST: In order to get more accurate snow depth readings, at 15:30 MST today, the SRRL BMS Snow Depth sensor was relocated from the north end of the radiometer tower to a new stand located 30 feet to the west of the radiometer tower. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]